# JEL classification: C11, C32, C55, E50

### v1

@@ 3 Application to Stock and Watson, (2012)
This section presents an empirical application based on the framework of Stock and Watson, (2012), who estimate a 30-variable vector autoregression (VAR) to assess the relevance of five macroeconomic shocks during the Great Financial Crisis. This empirical application serves the purpose to illustrate the effectiveness of our proposed algorithm to combine impact and shock sign restrictions in large a structural VAR. In line with their approach, we estimate a large-scale VAR model comprising the same set of variables, simultaneously identifying five structural shocks: an oil supply (news) shock, a monetary policy shock, a technology shock, a financial risk shock, and a government spending shock. [^3] Our empirical strategy closely follows Hou, (2024), who also estimate the same 30-variable VAR and identify the aforementioned structural shocks using a proxy SVAR approach. Moreover, we label our oil shock as an oil supply shock. In contrast to Stock and Watson, (2012), we extend the sample period to span from 1983Q1 to 2019Q3 and employ both impact sign and shock sign restrictions for the identification of structural shocks. [^4] Breitenlechner et al., (2024) note that this sample period captures a relatively stable macroeconomic environment, excluding the heightened oil price volatility of the 1970s as well as the COVID-19 pandemic and the geopolitical shock resulting from the Russian invasion of Ukraine. A detailed overview of the 30 variables included in the VAR is provided in Section Online Appendix A. Following the transformations employed in Chan et al., (2025), most variables are expressed in logarithms. To mitigate the risk of overfitting in such a high-dimensional setting, we impose horseshoe priors, as proposed in Korobilis, (2022) and Chan et al., (2025) and described above. The baseline model is estimated with a lag length of $p=4$, which balances model complexity with parsimony. Posterior inference is conducted using the described Gibbs sampler. Specifically, we perform 6,000 iterations, discarding the first 1,000 as burn-in and retaining every 10th of the remaining draws for inference.
@@ 3.1 Impact Sign and Shock Sign Restrictions
In this section, we outline and discuss the implementation of impact sign and shock sign restrictions. Particular emphasis is placed on the shock sign restrictions for 2001Q3 and 2008Q4, given their role in disentangling structural shocks. The remaining shock sign restrictions are deferred to Appendix Online Appendix B.
@@
Regarding the impact sign restrictions imposed on the $L$-matrix via $\ell_{i}^{L}<\bm{R}_{i}^{L}\bm{l}_{i}^{\prime}<\upsilon_{i}^{L}$ in Equation (7), we draw on a well-established literature on shock identification using sign restrictions. For the identification of a positive oil supply (news) shock, as presented in Table 1, we follow Banbura et al., (2023) and assume that such a shock leads to an increase in oil prices, as well as in both consumer and producer prices. [^5] Banbura et al., (2023) elaborate on the identification of oil supply and demand shocks, referring to Peersman and Van Robays, (2009), Kilian, (2008), Kilian, (2009), Kilian and Murphy, (2014), and Baumeister and Hamilton, (2019). Notably, our oil supply news is closely related to the shock identified by of Oil Supply News: Evidence from OPEC Announcements, (2021), where a positive shock implies the expectation of lower oil supply. In addition, we assume that a positive oil supply shock induces a monetary tightening, depresses stock prices, and worsens labor market conditions. Output and its components are left unrestricted. Concerning monetary policy shocks, we adopt the identification strategy of Chan et al., (2025), under which a tightening of monetary policy is associated with increases in interest rates and bond yields, declines in stock prices, and a contraction in economic activity. [^6] Extensive discussions on the identification of monetary policy shocks in VAR frameworks can be found in Sims, (1972), Sims, (1980), Sims, (1986), Faust, (1998), Canova and Pina, (2005), Canova and De Nicolo, (2002), Uhlig, (2005), and Rubio-Ramirez et al., (2010). The technology shock is identified following Chan et al., (2025), with the assumption that it increases economic activity, raises labor demand, and exerts downward pressure on prices. In response to such a shock, stock markets are expected to appreciate, while the policy rate declines in reaction to the deflationary environment. [^7] For additional discussion on the sign identification of technology shocks, see Dedola and Neri, (2007) and Peersman and Straub, (2009). To identify a financial risk shock, we assume a widening of the yield spread between corporate and ten year government bonds, a decline in stock market valuations, and a decrease in consumer prices, an approach consistent with Chan et al., (2025). [^8] See also Korobilis, (2022) and Furlanetto et al., (2019) for further discussion on the identification of financial shocks. Finally, a government spending shock is identified by assuming an increase in government expenditures, a reduction in unemployment, and a rise in (stock) prices. This identification strategy also aligns with Chan et al., (2025). [^9] See Laumer, (2020) and Forni et al., (2010) for detailed treatments of fiscal shock identification.
@@
As previously noted, we impose shock sign restrictions on the oil supply, financial risk, and government spending shocks in 2001Q3 and 2008Q4 through $\ell_{\tilde{t}}^{f}<\bm{R}_{\tilde{t}}^{f}\bm{f}_{\tilde{t}}<\upsilon_{\tilde{t}}^{f}$ in Equation (8), as these restrictions are essential for the separation of the respective shocks. In 2001Q3, the terrorist attacks on September 11th resulted in a sudden and pronounced increase in financial risk, reflecting heightened uncertainty across financial markets. The temporary closure of the New York Stock Exchange and the flight of investors toward safe-haven assets such as Treasury bonds substantiate this assessment. [^10] This interpretation is further supported by the Financial Stress Index; see Federal Reserve Bank of St. Louis, (2025). In terms of the oil shock, we assume it to be positive, as uncertainty surrounding oil supply rose significantly following the attacks, with market participants anticipating potential production cuts and hence, higher oil prices. Additionally, we assume a positive government spending shock in response to the enactment of the Emergency Supplemental Appropriations Act on September 18th, 2001.111111For further details, see Congress, (2001), Ramey and Zubairy, (2018), and Blanchard and Perotti, (2002). For the fourth quarter of 2008, we impose a negative oil supply news shock, a positive financial risk shock, and a negative government spending shock. The bankruptcy of Lehman Brothers in late September 2008 precipitated widespread turmoil in global financial markets, leading to a significant increase in financial uncertainty during the subsequent quarter. Consequently, we posit a positive financial risk shock.  [^12] See again Federal Reserve Bank of St. Louis, (2025) for empirical evidence. Regarding the oil supply shock, the financial crisis contributed to a downward revision of oil price expectations, thereby alleviating supply constraints.  [^13] In fact, the Organization of the Petroleum Exporting Countries (OPEC) held two meetings in 2008Q4 to address the decline in oil prices (Organization of the Petroleum Exporting Countries, , 2025). In terms of government spending, the election of President Obama resulted in a reduced expectation of government expenditure due to announcements of cuts in military spending.  [^14] Additionally, the American Recovery and Reinvestment Act, aimed at stimulating the economy, was announced by President Obama in January 2009 and signed into law in February 2009 (Congress, , 2009). As illustrated in Table 1, the narrative shock sign restrictions for both the financial risk and government spending shocks are critical for disentangling the two shocks. The additional shock sign restrictions further refine the identification process. It is noted that the monetary policy and the financial risk shock are not separated by means of the impact sign or shock sign restrictions. However, and as seen below in 3.2 and Online Appendix C the shock sign restrictions suffice. Further, we check the uniqueness of the sign patterns of the every retained shock impact matrix $L$ and the structural shocks $f_{t}$. [^15] The sign restrictions on shocks over distinct time periods may be interpreted as a more agnostic framework than impact sign restrictions, as noted in Antolín-Díaz and Rubio-Ramírez, (2018). Nonetheless, it is emphasized that sign restrictions must be grounded in a plausible narrative foundation.
@@
Impact Sign Restrictions Variable Oil Supply Monetary Policy Technology Financial Risk Government Spending GDP $-1$ $+1$ Consumption $-1$ $+1$ Investment $-1$ Gov. Spending $+1$ Hours $-1$ Real Compensation $+1$ $-1$ Output per Hour $+1$ Unit Labor Cost $+1$ $+1$ Industrial Production $-1$ Capacity Utilization Employment $-1$ $-1$ $+1$ Unemployment $+1$ $+1$ $-1$ $-1$ Housing Starts M&T Sales M&T Inventories PCE Index $+1$ $-1$ $-1$ $-1$ $+1$ GDP Deflator $+1$ $-1$ $-1$ $-1$ $+1$ CPI $+1$ $-1$ $-1$ $-1$ $+1$ PPI $+1$ $-1$ FED Funds Rates $+1$ $+1$ $-1$ 3-Month TB $+1$ 1-Year TB $+1$ 10-Year TB $+1$ BAA-GS10 Spread $-1$ $+1$ Monetary Base $-1$ M2 $-1$ S&P500 $-1$ $-1$ $+1$ $-1$ $+1$ Dow Jones $-1$ $-1$ $+1$ $-1$ $+1$ US Dollar Index $+1$ Oil Price $+1$ $+1$ Narrative Shock Sign Restrictions Oil Supply Monetary Policy Technology Financial Risk Government Spending $>0$ 1986Q3, 1988Q4 2001Q3, 2016Q4 2001Q3, 2008Q4 1990Q4, 2001Q3, 2001Q4, 2002Q1, 2002Q3, 2003Q1, 2006Q2, 2007Q4 $<0$ 2001Q4, 2008Q4 2014Q4 1998Q4, 1999Q4 1986Q3, 1988Q4, 1989Q4, 1991Q4, 2008Q4, 2011Q3, 2013Q1 Notes: This table summarizes the impact sign restrictions and narrative shock sign restrictions to identify the five structural shocks. In the upper table, $+1$ ($-1$) implies a positive (negative) on impact response of the respective variable in the left column. In the lower table, $>0$ ($<0$) restricts the respective shock in the column to be positive (negative).
@@ 3.2 Empirical Results
In this section, we present selected results from our empirical analysis. Figure 1 displays the impulse response functions (IRFs) of key macroeconomic and financial variables, including real GDP, industrial production, unemployment, the Consumer Price Index (CPI), and the S&P 500 index, in response to identified structural shocks. The oil supply news shock, which manifests as an increase in oil prices, leads to a contraction in economic activity, a decline in labor demand, and an increase in consumer prices. As expectations about future economic conditions deteriorate, equity markets also decline. These dynamics are consistent with the empirical findings in the literature (e.g., of Oil Supply News: Evidence from OPEC Announcements, (2021), Kilian and Park, (2009), Kilian, (2009), Hamilton, (2003), Ramey and Vine, (2011)). A contractionary monetary policy shock is found to reduce output, raise unemployment, and lower both prices and equity valuations (Christiano et al., (2005), Gertler and Karadi, (2015), Jarocinski and Karadi, (2020)). This response pattern suggests that the identified monetary policy shock is purged from any information effects or central bank signaling, thereby capturing a pure policy shock (see Jarocinski, (2022), Bauer and Swanson, 2023a , Bauer and Swanson, 2023b ). A positive technology shock increases both output and labor demand, while exerting downward pressure on prices due to productivity improvements. The stock market responds favorably. This is largely in line with existing evidence on the expansionary effects of technological innovations (Peersman and Straub, (2009)). A shock to financial risk has strongly contractionary effects: output and prices decline persistently, unemployment rises, and stock markets fall. These results corroborate prior findings on the macroeconomic consequences of financial stress (Gilchrist and ZakrajÅ¡ek, (2012), Caldara et al., (2016), Christiano et al., (2014), Jurado et al., (2015), Korobilis, (2022)). Finally, an increase in government spending yields an expansion in output, a reduction in unemployment, upward pressure on prices, and a short-run increase in equity prices. These responses are consistent with a substantial body of empirical work (Blanchard and Perotti, (2002), Ramey and Shapiro, (1998), Ramey, (2011), Auerbach and Gorodnichenko, (2012), Mountford and Uhlig, (2009), Laumer, (2020)). The complete set of impulse responses is provided in Section Online Appendix C of the Online Appendix.
@@
Table 2 presents the forecast error variance decomposition (FEVD) of GDP across various forecast horizons. The FEVD illustrates the proportion of the variance in GDP forecast errors that can be attributed to each of the identified structural shocks. Across all horizons, financial risk shocks emerge as the dominant driver of GDP fluctuations. Compared to the findings of Hou, (2024), our results suggest a greater role for oil supply shocks in driving the business cycle. This divergence may stem from the extended sample period employed in our empirical analysis. [^16] It should be noted that due to differences in identification strategies, our results are not directly comparable to those of Hou, (2024). Moreover, their sample is restricted to the period from 1987 to 2011, which is heavily influenced by the global financial crisis. In contrast, government spending shocks exhibit only a moderate contribution to the forecast error variance of GDP, while monetary policy and technology shocks account for a relatively smaller share of GDP fluctuations across all horizons. In the Online Appendix, Section Online Appendix C shows the historical decomposition of GDP corroborating the finding that financial risk shocks are the largest driver of the business cycle. Further, in Section Online Appendix C shows that if we shorten our sample and end in 2012 financial risk shocks take on an even more important role in driving the business cycle. [^17] Further results on the above-explained shortened sample can be obtained on request.

### v2

@@ 4 Application to Stock and Watson, (2012)
This section presents an empirical application based on the framework of Stock and Watson, (2012), who estimate a 30-variable vector autoregression (VAR) to assess the relevance of five macroeconomic shocks during the Great Financial Crisis. This empirical application serves the purpose to illustrate the effectiveness of our proposed algorithm to combine impact and shock sign restrictions in large a structural VAR. [^10] Hou, (2024) also estimates the same 30-variable VAR and identifies the structural shocks using multiple instruments in a proxy SVAR. In line with their approach, we estimate a large-scale VAR model comprising the same set of variables, simultaneously identifying five structural shocks: an oil supply (news) shock, a monetary policy shock, a technology shock, a financial risk shock, and a government spending shock. Differing from Stock and Watson, (2012), we employ the above-outlined structural VAR with factor error structure. Further, we extend the sample period to span from 1983Q1 to 2019Q4 and employ both impact sign and shock sign restrictions for the identification of structural shocks. [^11] Breitenlechner et al., (2024) note that this sample period captures a relatively stable macroeconomic environment, excluding the heightened oil price volatility of the 1970s as well as the COVID-19 pandemic and the geopolitical shock resulting from the Russian invasion of Ukraine. A detailed overview of the 30 variables included in the VAR is provided in Section Online Appendix A. Following the transformations employed in Chan et al., (2025), most variables are expressed in logarithms. Given the quarterly frequency of the data, the model is estimated using a lag length of $p=4$. Posterior inference is conducted using the described Gibbs sampler. Specifically, we perform 6,000 iterations, discarding the first 1,000 as burn-in and retaining every 10th of the remaining draws for inference.
@@ 4.1 Impact Sign and Shock Sign Restrictions
In this section, we outline and discuss the implementation of impact sign and shock sign restrictions. Particular emphasis is placed on the shock sign restrictions for 2001Q3 and 2008Q4, given their role in disentangling structural shocks. The reasoning of the remaining shock sign restrictions are deferred to Appendix Online Appendix B.
@@
Regarding the impact sign restrictions imposed on the $L$-matrix via $\ell_{i}^{L}<\bm{R}_{i}^{L}\bm{l}_{i}^{\prime}<\upsilon_{i}^{L}$ in Equation (7), we draw on a well-established literature on shock identification using sign restrictions. For the identification of a positive oil supply (news) shock, as presented in Table 2, we follow Banbura et al., (2023) and assume that such a shock leads to an increase in oil prices, as well as in both consumer and producer prices. [^12] Banbura et al., (2023) elaborate on the identification of oil supply and demand shocks, referring to Peersman and Van Robays, (2009), Kilian, (2008), Kilian, (2009), Kilian and Murphy, (2014), and Baumeister and Hamilton, (2019). Notably, our oil supply news is closely related to the shock identified by Känzig, (2021), where a positive shock implies the expectation of lower oil supply. In addition, we assume that a positive oil supply shock induces a monetary tightening, depresses stock prices, and worsens labor market conditions. Output and its components are left unrestricted. Concerning monetary policy shocks, we adopt the identification strategy of Chan et al., (2025), under which a tightening of monetary policy is associated with increases in interest rates and bond yields, declines in stock prices, and a contraction in economic activity. [^13] Extensive discussions on the identification of monetary policy shocks in VAR frameworks can be found in Sims, (1972), Sims, (1980), Sims, (1986), Faust, (1998), Canova and Pina, (2005), Canova and De Nicolo, (2002), Uhlig, (2005), and Rubio-Ramirez et al., (2010). The technology shock is identified following Chan et al., (2025), with the assumption that it increases economic activity, raises labor demand, and exerts downward pressure on prices. In response to such a shock, stock markets are expected to appreciate, while the policy rate declines in reaction to the deflationary environment. [^14] For additional discussion on the sign identification of technology shocks, see Dedola and Neri, (2007) and Peersman and Straub, (2009). To identify a financial risk shock, we assume a widening of the yield spread between corporate and ten year government bonds, a decline in stock market valuations, and a decrease in consumer prices, an approach consistent with Chan et al., (2025). [^15] See also Korobilis, (2022) and Furlanetto et al., (2019) for further discussion on the identification of financial shocks. Finally, a government spending shock is identified by assuming an increase in government expenditures, a reduction in unemployment, and a rise in (stock) prices. This identification strategy also aligns with Chan et al., (2025). [^16] See Laumer, (2020) and Forni et al., (2010) for detailed treatments of fiscal shock identification.
@@
As previously noted, we impose shock sign restrictions on the oil supply, financial risk, and government spending shocks in 2001Q3 and 2008Q4 through $\ell_{\tilde{t}}^{f}<\bm{R}_{\tilde{t}}^{f}\bm{f}_{\tilde{t}}<\upsilon_{\tilde{t}}^{f}$ in Equation (8), as these restrictions are essential for the separation of the respective shocks. In 2001Q3, the terrorist attacks on September 11th resulted in a sudden and pronounced increase in financial risk, reflecting heightened uncertainty across financial markets. The temporary closure of the New York Stock Exchange and the flight of investors toward safe-haven assets such as Treasury bonds substantiate this assessment. [^17] This interpretation is further supported by the Financial Stress Index; see Federal Reserve Bank of St. Louis, (2025). In terms of the oil shock, we assume it to be positive, as uncertainty surrounding oil supply rose significantly following the attacks, with market participants anticipating potential production cuts and hence, higher oil prices. Additionally, we assume a positive government spending shock in response to the enactment of the Emergency Supplemental Appropriations Act on September 18th, 2001.181818For further details, see Congress, (2001), Ramey and Zubairy, (2018), and Blanchard and Perotti, (2002). For the fourth quarter of 2008, we impose a negative oil supply news shock, a positive financial risk shock, and a negative government spending shock. The bankruptcy of Lehman Brothers in late September 2008 precipitated widespread turmoil in global financial markets, leading to a significant increase in financial uncertainty during the subsequent quarter. Consequently, we posit a positive financial risk shock.  [^19] See again Federal Reserve Bank of St. Louis, (2025) for empirical evidence. Regarding the oil supply shock, the financial crisis contributed to a downward revision of oil price expectations, thereby alleviating supply constraints.  [^20] In fact, the Organization of the Petroleum Exporting Countries (OPEC) held two meetings in 2008Q4 to address the decline in oil prices (Organization of the Petroleum Exporting Countries, , 2025). In terms of government spending, the election of President Obama resulted in a reduced expectation of government expenditure due to announcements of cuts in military spending.  [^21] Additionally, the American Recovery and Reinvestment Act, aimed at stimulating the economy, was announced by President Obama in January 2009 and signed into law in February 2009 (Congress, , 2009). As illustrated in Table 2, the narrative shock sign restrictions for both the financial risk and government spending shocks are critical for disentangling the two shocks. The additional shock sign restrictions further refine the identification process. It is noted that, for the monetary policy and financial risk shocks, in principle, it would be possible for the two shocks to exhibit the same sign patterns. However, given our set of restrictions, the data do not support the same sign patterns for both shocks, and thus our restrictions are sufficient to disentangle the shocks. In particular, we verify the uniqueness of the sign patterns for every posterior draw of the impact matrix $L$ and the structural shocks $f_{t}$.
@@
Impact Sign Restrictions Variable Oil Supply Monetary Policy Technology Financial Risk Government Spending GDP 1-1- 1 +1+1+ 1 Consumption 1-1- 1 +1+1+ 1 Investment 1-1- 1 Gov. Spending +1+1+ 1 Hours 1-1- 1 Real Compensation +1+1+ 1 1-1- 1 Output per Hour +1+1+ 1 Unit Labor Cost +1+1+ 1 +1+1+ 1 Industrial Production 1-1- 1 Capacity Utilization Employment 1-1- 1 1-1- 1 +1+1+ 1 Unemployment +1+1+ 1 +1+1+ 1 1-1- 1 1-1- 1 Housing Starts M&T Sales M&T Inventories PCE Index +1+1+ 1 1-1- 1 1-1- 1 1-1- 1 +1+1+ 1 GDP Deflator +1+1+ 1 1-1- 1 1-1- 1 1-1- 1 +1+1+ 1 CPI +1+1+ 1 1-1- 1 1-1- 1 1-1- 1 +1+1+ 1 PPI +1+1+ 1 1-1- 1 FED Funds Rates +1+1+ 1 +1+1+ 1 1-1- 1 3-Month TB +1+1+ 1 1-Year TB +1+1+ 1 10-Year TB +1+1+ 1 BAA-GS10 Spread 1-1- 1 +1+1+ 1 Monetary Base 1-1- 1 M2 1-1- 1 S&P500 1-1- 1 1-1- 1 +1+1+ 1 1-1- 1 +1+1+ 1 Dow Jones 1-1- 1 1-1- 1 +1+1+ 1 1-1- 1 +1+1+ 1 US Dollar Index +1+1+ 1 Oil Price +1+1+ 1 +1+1+ 1 Narrative Shock Sign Restrictions Oil Supply Monetary Policy Technology Financial Risk Government Spending >0>0> 0 1986Q3, 1988Q4 2001Q3, 2016Q4 2001Q3, 2008Q4 1990Q4, 2001Q3, 2001Q4, 2002Q1, 2002Q3, 2003Q1, 2006Q2, 2007Q4 <0<0< 0 2001Q4, 2008Q4 2014Q4 1998Q4, 1999Q4 1986Q3, 1988Q4, 1989Q4, 1991Q4, 2008Q4, 2011Q3, 2013Q1 Notes: This table summarizes the impact sign restrictions and narrative shock sign restrictions to identify the five structural shocks. In the upper table, +1+1+ 1 (1-1- 1) implies a positive (negative) on impact response of the respective variable in the left column. In the lower table, >0>0> 0 (<0<0< 0) restricts the respective shock in the column to be positive (negative).
@@ 4.2 Empirical Results
In this section, we present selected results from our empirical analysis. Figure 1 displays the impulse response functions (IRFs) of key macroeconomic and financial variables, including real GDP, industrial production, unemployment, the Consumer Price Index (CPI), and the S&P 500 index, in response to identified structural shocks. The oil supply news shock, which manifests as an increase in oil prices, leads to a contraction in economic activity, a decline in labor demand, and an increase in consumer prices. As expectations about future economic conditions deteriorate, equity markets also decline. These dynamics are consistent with the empirical findings in the literature (e.g., Känzig, (2021), Kilian and Park, (2009), Kilian, (2009), Hamilton, (2003), Ramey and Vine, (2011)). A contractionary monetary policy shock is found to reduce output, raise unemployment, and lower both prices and equity valuations (Christiano et al., (2005), Gertler and Karadi, (2015), Jarocinski and Karadi, (2020)). This response pattern suggests that the identified monetary policy shock is purged from any information effects or central bank signaling, thereby capturing a pure policy shock (see Jarocinski, (2022), Bauer and Swanson, 2023a , Bauer and Swanson, 2023b ). A positive technology shock increases both output and labor demand, while exerting downward pressure on prices due to productivity improvements. The stock market responds favorably. This is largely in line with existing evidence on the expansionary effects of technological innovations (Peersman and Straub, (2009)). A shock to financial risk has strongly contractionary effects: output and prices decline persistently, unemployment rises, and stock markets fall. These results corroborate prior findings on the macroeconomic consequences of financial stress (Gilchrist and Zakrajšek, (2012), Caldara et al., (2016), Christiano et al., (2014), Jurado et al., (2015), Korobilis, (2022)). Finally, an increase in government spending yields an expansion in output, a reduction in unemployment, upward pressure on prices, and a short-run increase in equity prices. These responses are consistent with a substantial body of empirical work (Blanchard and Perotti, (2002), Ramey and Shapiro, (1998), Ramey, (2011), Auerbach and Gorodnichenko, (2012), Mountford and Uhlig, (2009), Laumer, (2020)). The complete set of impulse responses is provided in Section Online Appendix C of the Online Appendix.
@@
Table 3 presents the forecast error variance decomposition (FEVD) of GDP across various forecast horizons. The FEVD illustrates the proportion of the variance in GDP forecast errors that can be attributed to each of the identified structural shocks. Across all horizons, financial risk shocks emerge as the dominant driver of GDP fluctuations. In contrast, government spending shocks exhibit only a moderate contribution to the forecast error variance of GDP, while monetary policy and technology shocks account for a relatively smaller share of GDP fluctuations across all horizons. Our empirical results are in line with Hou, (2024) who estimates the same 30-variable VAR and proxy-identifies the 5 structural shocks. In the Online Appendix, Section Online Appendix C shows the historical decomposition of GDP corroborating the finding that financial risk shocks are the largest driver of the business cycle. Further, in Section Online Appendix C we show that if the sample is shortened to 2012 financial risk shocks take on an even more important role in driving the business cycle. [^22] Further results on the above-explained shortened sample can be obtained on request. It is noted that our results from the shortened sample are even more similar to Hou, (2024) as their baseline sample ends in 2012.